ROTATIONAL AND HYPERFINE ANALYSIS OF THE a^{1}Δ_{g}←X^{3}Σ_{g}^{−} BAND OF ^{17}O-CONTAINING ISO-TOPOLOGUES OF OXYGEN MEASURED BY CRDS AT ROOM AND LIQUID NITROGEN TEMPERA-TURES

O. M. LESHCHISHINA, S. KASSI, Université de Grenoble, CNRS UMR 5588, LIPHY, 38041 Grenoble, France; \underline{I. E. GORDON}, Harvard-Smithsonian Center for Astrophysics, Atomic and Molecular Physics Division, Cambridge MA 02138-1516, USA; S. YU, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA 91109, USA; A. CAMPARGUE, Université de Grenoble, CNRS UMR 5588, LIPHY, 38041 Grenoble, France.

The a^{1}Δ_{g}←X^{3}Σ_{g}^{−} system of the ^{16}O^{17}O, ^{17}O^{18}O and ^{17}O_{2} isotopologues of oxygen was studied by high sensitivity CW-Cavity Ring Down Spectroscopy. The spectra of a ^{17}O highly enriched sample were recorded at room temperature between 7640 and 7917 cm^{−1} and at liquid nitrogen temperature in the 7876-7893 cm^{−1} region. The magnetic dipole (0-0) band was observed for all three ^{17}O-containing isotopologues. At liquid nitrogen temperature some of the transitions were observed with partially resolved hyperfine splitting due to the ^{17}O nuclear spin, allowing determination of the hyperfine constants. The electric quadrupole (0-0) band and the (1-1) magnetic dipole hot band were also observed for the ^{16}O^{17}O and ^{17}O_{2} species. The rotational and hyperfine spectroscopic parameters of the X^{3}Σ_{g}^{−} and a^{1}Δ_{g} states of the three studied isotopologues were derived from global fit of the measured line positions and microwave and Raman measurements available in the literature. The rotational constants of the a^{1}Δ_{g} (v=0, 1) states of ^{17}O_{2} are determined for the first time.